### v1

@@ 2.4 Avoiding Barren Plateaus In Complex Models
$\frac{\partial\langle O\rangle_{\rho,\mathbf{\theta}}}{\partial\theta_{\mu}}=\frac{1}{4}(X^{(\mu)}_{1}+X^{(\mu)}_{2}+X^{(\mu)}_{3}+X^{(\mu)}_{4}).$ (56)
A model is said to contain barren plateaus if $\forall\theta_{\mu}\in\mathbf{\theta}$ the following condition holds Larocca_2022
where $\text{Var}_{\mathbf{\theta}}$ is the variance calculated over a uniform distribution of parameters $\mathbf{\theta}$. If we calculate this gradient variance term for the quantum ResNet example, we see that
@@
$\frac{1}{4}\big{(}\text{Var}_{\mathbf{\theta}}(X_{1})+\text{Var}_{\mathbf{\theta}}(X_{2})+\text{Var}_{\mathbf{\theta}}(X_{3})+\text{Var}_{\mathbf{\theta}}(X_{4})$
$\negthinspace\negthinspace+\negthinspace\text{Cov}_{\mathbf{\theta}}(X_{1},X_{2})+\text{Cov}_{\mathbf{\theta}}(X_{1},X_{3})+\text{Cov}_{\mathbf{\theta}}(X_{1},X_{4})$
$\negthinspace\negthinspace+\negthinspace\text{Cov}_{\mathbf{\theta}}(X_{2},X_{3})+\text{Cov}_{\mathbf{\theta}}(X_{2},X_{4})+\text{Cov}_{\mathbf{\theta}}(X_{3},X_{4})\big{)},$ (58)
where $\text{Cov}_{\mathbf{\theta}}$ is the covariance between two variables calculated over a uniform distribution of parameters $\mathbf{\theta}$ and we have dropped the superscripts of $\mu$ on the $X$ terms for clarity of notation. There are many documented variational circuits that do not contain barren plateaus Schatzki_2024 ; wiersema2023classification ; west2024provably ; diaz2023showcasing . Hence, selecting $W_{2}(\mathbf{\theta}_{2}),\rho,O$ to lie within any of these frameworks will lead to a situation in which the term $\text{Var}_{\mathbf{\theta}}(X_{1})$ exhibits polynomial scaling $\text{Var}_{\mathbf{\theta}}(X_{1})=\mathcal{O}(\frac{1}{\text{poly(n)}})$. If this is the case, then it is possible that the gradient variance for the entire model could scale as $\text{Var}_{\mathbf{\theta}}(\frac{\partial\langle O\rangle_{\rho,{\mathbf{\theta}}}}{\partial\theta_{\mu}})=\mathcal{O}(\frac{1}{\text{poly(n)}})$, and hence barren plateaus will be avoided due to this beneficially scaling term under certain conditions: Firstly, the number of layers must scale as $L=\mathcal{O}(\log(n))$ in the general case, otherwise the beneficially scaling term may be exponentially suppressed. This requirement can be relaxed to $L=\mathcal{O}(\text{poly}(n))$ if considering special cases where less residual connections are used overall, as in the example above by setting some $\beta=1$ or $\beta=0$, or as demonstrated in the special case outlined in Appendix E. Secondly, it would also require that the covariance terms do not cancel out the beneficially scaling term.
The unique feature of this quantum ResNet implementation lies in its ability to sidestep barren plateaus while incorporating complex variational ansätze. Specifically, if the conditions for avoiding barren plateaus are met when considering only the $W_{2}(\mathbf{\theta}_{2})$ variational operator, then the $W_{1}(\mathbf{\theta}_{1})$ layer can be freely chosen to be any variational ansatz, even those known to exhibit barren plateaus, and the overall model could still avoid barren plateaus. This flexibility allows $W_{1}(\mathbf{\theta}_{1})$ to be chosen to be a circuit that is hard to simulate classically, ensuring that the overall model remains difficult to simulate. In other words, you may only need one layer that does not exhibit barren plateaus to avoid them completely, and you only need one layer that is hard to simulate classically to ensure the entire model is hard to simulate overall.
This approach provides a novel method for constructing VQC models that may avoid the trainability-classical simulability issue prevalent in the current literature; it has been reported that a wide class of VQCs that do not exhibit barren plateaus are classically simulatable cerezo2023does , limiting the potential for quantum advantage. For example, it has been shown that for certain classes of VQC models, a likely requirement for trainability (the absence of barren plateaus) is that the dynamical Lie Algebra (DLA) dimension of the VQC is polynomial in size fontana2023adjoint ; ragone2023unified . Simultaneously, many VQC models with a polynomially sized DLA have also been shown to be classically simulatable goh2023liealgebraic , bringing into question their potential for quantum advantage. Beyond seeking quantum advantage for its own sake, it has also recently been demonstrated that polynomial sized DLA VQC models of this type are vulnerable to “weak privacy breaches” in which expectation value snapshots of quantum-encoded confidential data can be retrieved from the training gradients shared between nodes in a distributed learning setting; allowing potential attack agents to learn information about confidential input data heredge2024prospects . This adds further urgency to find more complex quantum models that can still avoid the trainability issues associated with barren plateaus. However, by integrating a hard-to-simulate $W_{1}(\mathbf{\theta}_{1})$ (e.g., exponential size DLA) with a carefully designed trainable $W_{2}(\mathbf{\theta}_{2})$ (e.g., polynomial DLA), our model has the potential to remain non-classically simulatable while also remaining trainable. This potentially paves the way for the development of VQC models that leverage quantum complexity while maintaining efficient trainability. Furthermore, as the quantum ResNet construction allows for non-unitary operations that result in loss functions with parts deriving from both short and deep circuits simultaneously this means that many standard methods and assumptions used for analysing VQC circuits may not apply, leaving ample opportunity for further research.
Note that the presence of non-unitary evolution terms, such as $X_{3}$ and $X_{4}$, in the loss function gradient also introduces additional considerations regarding simulatability. As in general, these terms will not even be calculable on a quantum device using strictly unitary operations. The impact of these terms on simulatability is left as a subject for further research.
3 Average Pooling Layers
3.1 Quantum Average Pooling Implementation

### v2

@@ II.4.1 Avoiding Barren Plateaus
$\langle O\rangle_{\rho,\mathbf{\theta}}$ $=\frac{1}{\Omega^{\prime 2}}\Big{(}(1-\beta)^{2}\text{Tr}(W_{2}(\mathbf{\theta}_{2})\rho W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$
$+\beta^{2}\text{Tr}(W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})\rho W_{1}(\mathbf{\theta}_{1})^{\dagger}W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$
$+\beta(1-\beta)\text{Tr}(W_{2}(\mathbf{\theta}_{2})\tilde{\rho}_{\mathbf{\theta}_{1}}W_{2}(\mathbf{\theta}_{2})^{\dagger}O),$ (53)
where $\tilde{\rho}_{\mathbf{\theta}_{1}}=\rho W_{1}(\mathbf{\theta}_{1})^{\dagger}+W_{1}(\mathbf{\theta}_{1})\rho$. It is now possible to identify three distinct terms in the loss function.
Deep - $L_{\text{BP}}=\text{Tr}(W_{2}(\mathbf{\theta}_{2})W_{1}(\mathbf{\theta}_{1})\rho W_{1}(\mathbf{\theta}_{1})^{\dagger}W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$
This component is identical to the unitary VQC loss function if no residual component had been implemented. It corresponds to a deep circuit consisting of $W_{1}(\mathbf{\theta}_{1})$ followed by $W_{2}(\mathbf{\theta}_{2})$ in series, which we will assume exhibits barren plateaus.
Shallow - $L_{\text{No-BP}}=\text{Tr}(W_{2}(\mathbf{\theta}_{2})\rho W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$
This component is identical to the unitary VQC loss function if the variational circuit consisted of $W_{2}(\mathbf{\theta}_{2})$ only. We will assume that this model does not exhibit barren plateaus, as has been shown to be the case for many variational circuits Schatzki et al. (2024); Wiersema et al. (2023); West et al. (2024); Diaz et al. (2023), but may be vulnerable to being classically simulated Cerezo et al. (2023).
Non-unitary - $L_{\tilde{\rho}}=\text{Tr}(W_{2}(\mathbf{\theta}_{2})\tilde{\rho}_{\mathbf{\theta}_{1}}W_{2}(\mathbf{\theta}_{2})^{\dagger}O)$
This component contains the non-unitary element of the loss function. It appears similar to the $L_{\text{No-BP}}$ term with the key difference that instead of $\rho$ being unitary evolved we have $\tilde{\rho}_{\mathbf{\theta}_{1}}=\rho W_{1}(\mathbf{\theta}_{1})^{\dagger}+W_{1}(\mathbf{\theta}_{1})\rho$. While $\tilde{\rho}_{\mathbf{\theta}_{1}}$ is Hermitian, it does not necessarily have a trace equal to 1 nor is it positive semidefinite in general.
Figure 5: Visualisation of the single residual connection $L=1$ quantum ResNet loss function $\langle O\rangle_{\rho,\mathbf{\theta}}$ when decomposed into the individual terms corresponding to the shallow component $L_{\text{No-BP}}$, deep component $L_{\text{BP}}$ and non-unitary component $L_{\tilde{\rho}}$. In the non-unitary component the arrow corresponds to the transformation $\rho\rightarrow\tilde{\rho}=\rho W_{1}^{\dagger}+W_{1}\rho$.
Setting $\beta=\frac{1}{2}$ and absorbing this factor along with the normalisation constant into the loss definition we can write
$\langle O\rangle_{\rho,\mathbf{\theta}}=L_{\text{No-BP}}+L_{\text{BP}}+L_{\tilde{\rho}}.$ (54)
A visual interpretation of this resulting quantum ResNet model is shown in Figure 5. We can therefore find the gradient of this loss function with respect to parameters $\theta_{\mu}\in\mathbf{\theta}$ as
$\frac{\partial\langle O\rangle_{\rho,\mathbf{\theta}}}{\partial\theta_{\mu}}=\frac{\partial L_{\text{No-BP}}}{\partial\theta_{\mu}}+\frac{\partial L_{\text{BP}}}{\partial\theta_{\mu}}+\frac{\partial L_{\tilde{\rho}}}{\partial\theta_{\mu}}.$ (55)
The resulting gradient variance term can then be written as
@@
$\text{Var}_{\mathbf{\theta}}(\frac{\partial L_{\text{No-BP}}}{\partial\theta_{\mu}})+\text{Var}_{\mathbf{\theta}}(\frac{\partial L_{\text{BP}}}{\partial\theta_{\mu}})+\text{Var}_{\mathbf{\theta}}(\frac{\partial L_{\tilde{\rho}}}{\partial\theta_{\mu}})$
$\negthinspace\negthinspace+\negthinspace\text{Cov}_{\mathbf{\theta}}(\frac{\partial L_{\text{No-BP}}}{\partial\theta_{\mu}},\frac{\partial L_{\text{BP}}}{\partial\theta_{\mu}})+\text{Cov}_{\mathbf{\theta}}(\frac{\partial L_{\text{No-BP}}}{\partial\theta_{\mu}},\frac{\partial L_{\tilde{\rho}}}{\partial\theta_{\mu}})$
$\negthinspace\negthinspace+\negthinspace\text{Cov}_{\mathbf{\theta}}(\frac{\partial L_{\text{BP}}}{\partial\theta_{\mu}},\frac{\partial L_{\tilde{\rho}}}{\partial\theta_{\mu}}).$ (56)
By construction we know that $L_{\text{No-BP}}$ does not lead to barren plateaus and we can assume that $\text{Var}_{\mathbf{\theta}}(\frac{\partial L_{\text{No-BP}}}{\partial\theta_{\mu}})\sim\frac{1}{\text{poly}(n)}$ does not decay exponentially. Hence the overall quantum ResNet model has gradient variance $\text{Var}_{\mathbf{\theta}}\left(\frac{\partial\langle O\rangle_{\rho,{\mathbf{\theta}}}}{\partial\theta_{\mu}}\right)\sim\frac{1}{\text{poly}(n)}$ and therefore does not exhibit barren plateaus. The shallow depth component has mitigated barren plateaus in the overall model. This is demonstrated experimentally in Figure 6 which shows that the gradient variance $\text{Var}_{\mathbf{\theta}}\left(\frac{\partial\langle O\rangle_{\rho,{\mathbf{\theta}}}}{\partial\theta_{\mu}}\right)$ for the overall quantum ResNet model decays sub-exponentially.
Figure 6: The logarithm of the gradient variance as the number of qubits is increased. (Quantum Residual Connections) The model in which one residual layer is implemented between the input and output of the $W_{1}$ portion of the BP ansatz which consists of $W_{2}W_{1}$. (No Residual Connections) The model with no residual connections is an ansatz which is known to exhibit barren plateaus and corresponds to the operator $W_{2}W_{1}$. In this example $W_{1}$ was generated by $[XY,YX,YZ]$ leading to an exponential DLA dimension Wiersema et al. (2023) while $W_{2}$ was generated by $[XY]$ leading to a polynomial DLA dimension Wiersema et al. (2023). Both $W_{1}$ and $W_{2}$ consisted of 25 repeated sublayers of parameterised gates. Each sublayer consisted of parameterised two-qubit gates that acted on all adjacent qubits of the form $\prod_{k=1}\big{(}(\bigotimes_{j=\text{odd}}e^{i\theta_{j}H^{(k)}_{j,j+1}})(\bigotimes_{j^{\prime}=\text{even}}e^{i\theta_{j^{\prime}}H^{(k)}_{j^{\prime},j^{\prime}+1}})\big{)}$ where $H^{(k)}_{j,j+1}$ is the $k$-th element in the specified set of generators which acts on the qubits indexed by $j$ and $j+1$. Variational parameters $\theta\in\mathbf{\theta}$ were sampled over the range $\theta\in[0,2\pi]$ and 1000 samples were taken in total. The initial state $\rho_{0}=\lvert 0\rangle^{\otimes n}\langle 0\rvert^{\otimes n}$ and the final observable was $O=Y\otimes Y\otimes\mathbb{I}^{n-2}$.
II.4.2 Classical Simulation
It has recently been reported that a broad class of VQC models that avoid barren plateaus are also classically simulatable Cerezo et al. (2023), limiting their potential for quantum advantage. While quantum residual connections offer a means to circumvent barren plateaus, they may still lead to scenarios where the model is well-approximated by the shallow component of the loss function $L_{\text{No-BP}}$, which may be classical simulatable Cerezo et al. (2023). By construction we expect the $L_{\text{BP}}$ to exponentially vanish, and if the same happens to the $L_{\tilde{\rho}}$ term then this implies that the entire quantum ResNet model could be approximated efficiently on a classical computer within a given error. This raises the question: Are quantum ResNet models that avoid barren plateaus classically simulatable in general? If we take the assumption that $L_{\text{BP}}$ exponentially vanishes and $L_{\text{No-BP}}$ is classically simulatable, this means the answer depends on the characteristics of the non-unitary term $L_{\tilde{\rho}}$. If the $L_{\tilde{\rho}}$ term does not decay exponentially and cannot be classically simulated, then this condition could lead to a situation in which a model with at least one residual connection could be difficult to simulate, and simultaneously avoid barren plateaus. This non-unitary term, has not yet been characterised in general and therefore highlights a possible search space for future research, however, we note that this may not necessarily be a common situation. We show in Figure 7 that for the particular quantum ResNet architecture we consider in our numerics, that $L_{\tilde{\rho}}$ exponentially decays with an increasing number of qubits, which suggests that this specific architecture may be vulnerable to classical simulation. However, we only consider a limited example and note that there is a large space which remains unknown and is open to further work. Generalised quantum operations of the form $\sum_{i}\alpha_{i}U_{i}$, which are implemented by the LCU procedure, have been shown to form a convex set and the extreme points of this set are the unitary operations Gudder (2007). This means that the space of operators that can be considered is much larger than that which has previously been studied in the literature, and a full characterisation of this general space is beyond the scope of this current work.
Figure 7: The logarithm of the average magnitude of the non-unitary term $L_{\tilde{\rho}}$. The average was performed over 1000 initialisation. Data is from the same experiment specified in Figure 6. In this particular setting the non-unitary contribution vanishes exponentially, suggesting it may be possible to approximate the model by $L_{\text{BP}}+L_{\text{No-BP}}$. If the $L_{\text{No-BP}}$ loss vanishes exponentially too, then the entire quantum ResNet could be approximated by $L_{\text{BP}}$.
II.4.3 Quantum ResNet as an Ensemble
In the following portion we use the term layer to refer to quantum ResNet layers, such that $L$ corresponds to the number residual connections and hence the number of controlled unitaries in the model. Each individual controlled unitary $W_{l}$ is a variational ansatz, which in many cases is composed of repeated layers of a certain gate architecture; to avoid confusion, we will refer to the repeated layers within these unitary $W_{l}$ terms as sublayers. We will now denote the unitary in the $l$-th residual network layer as $W_{l}^{(\lambda)}$, where $\lambda$ specifies the number of repeated sublayers in the ansatz corresponding to that particular unitary.
Previous work on the characterisation of barren plateaus makes assumptions that the circuits involved are sufficiently deep Ragone et al. (2024); Fontana et al. (2023). As the quantum ResNet allows layer skipping, this means that there may always be shallow depth circuit components in the final loss function, which will not satisfy this assumption. We now consider a deep circuit composed of many repeated sublayers of the same structure of variational gates. We show that it is possible with $L$ layers of quantum ResNet to effectively implement an ensemble of $2^{L}$ different ansätze where the number of sublayers of the individual ansatz terms in the ensemble ranges from $1$ to $2^{L}$.
The uniform ensemble quantum ResNet architecture is created using $L$ quantum residual layers by ensuring the number of repeated sublayers $\lambda$ in the $l$-th unitary $W_{l}^{(\lambda)}$ is given by $\lambda=2^{l-1}$ for $1\leq l\leq L$. To prevent a constant term in the loss function arising from the identity, we also initially apply a unitary operator consisting of a single sublayer $\lambda=1$, without using a residual connection, which we denote as $W_{0}^{(1)}$. This corresponds to applying the operator
$(\mathbb{I}+W_{L}^{(2^{L-1})})...(\mathbb{I}+W_{3}^{(4)})(\mathbb{I}+W_{2}^{(2)})(\mathbb{I}+W_{1}^{(1)})W_{0}^{(1)},$ (57)
and normalising the resultant state by a factor $\frac{1}{\Omega^{\prime}}$. This term expands out to give an operation consisting of a linear combination of $2^{L}$ terms. When unitaries multiply to create these terms, the number of sublayers add together. It is therefore clear that this procedure results in a operator
$\sum_{\lambda^{\prime}=1}^{2^{L}}V^{(\lambda^{\prime})},$ (58)
where $V^{(\lambda^{\prime})}$ is a unitary corresponding to
$V^{(\lambda^{\prime})}=\Big{(}\prod_{k\in S_{\lambda^{\prime}}}W_{k}^{(2^{k-1})}\Big{)}W_{0}^{(1)}.$ (59)
Where $S_{\lambda^{\prime}}\subseteq[L,1]$ is the set that indexes the $W_{k}^{(2^{k-1})}$ terms that multiply together to give a total number of sublayers $\lambda^{\prime}=1+\sum_{k\in S_{\lambda^{\prime}}}2^{k-1}$. The state is also normalised by a factor $\frac{1}{\Omega^{\prime\prime}}$ after the operation is applied. In total, we have $2^{L}$ terms where the number of sublayers in each terms is in the range $\lambda^{\prime}\in[1,2^{L}]$. This will lead to a loss function with $2^{L}$ unitary VQC model loss functions, covering sublayer depths from $1$ to $2^{L}$, in addition to non-unitary cross terms.
Many investigations into barren plateaus specifically require circuits to be of a certain depth, for example, requiring sufficient depth to form an approximate design over a Lie group Ragone et al. (2024). While bounds have been found for this depth for certain circuits, the quantum ResNet would allow an agnostic approach where many different layers can be trialled in the same model run and where shallow circuit components could guarantee the overall model does not exhibit barren plateaus. We see that for $L$ quantum ResNet layers one can implement $2^{L}$ different unitary terms in an ensemble, in addition to further non-unitary cross terms.
Loss function concentration implies loss function gradient concentration, and hence this metric can be similarly used to discuss the presence of barren plateaus Arrasmith et al. (2022). An ensemble sum of all terms with differing numbers of sublayers may be able to avoid barren plateaus due to the presence of shallower depth components within the average ensemble, as shown in Figure 8. This reaffirms previous results showing quantum ensemble methods can avoid barren plateaus Friedrich and Maziero (2024). However, the quantum ResNet will also provide additional sources of variance from the non-unitary terms in the loss function, although a detailed characterisation of these terms is beyond the scope of this work.
Note that as the number of quantum ResNet layers $L$ increases, the number of terms grows exponentially, meaning that any individual term will have a weighting that decays exponentially due to the normalisation of the state after the operation is applied. Therefore, if $L$ is increased, one would need to ensure a sufficient amount of non-exponentially concentrating terms survive to still avoid barren plateaus. By initialising the ancilla qubits, one can vary the weightings of particular terms in the ensemble. This flexibility could allow the creation of models that can be pushed close to the boundary of barren plateaus but still remain in the trainable region. Similarly to the arguments made for the previous model, if the non-unitary terms vanish, then we would expect the loss function to be on average well approximated by the shallow non-vanishing terms which do not exhibit barren plateaus individually (excluding rare potential configurations in which deep terms still provide non-zero contributions). If these shallow terms turn out to be classically simulatable Cerezo et al. (2023) then this means that the overall quantum ResNet will be classically simulatable to within some given error. Efforts to avoid this situation should focus on finding settings in which the non-unitary terms do not exponentially vanish; it remains an open question whether this is possible.
Figure 8: The logarithm of the variance of the loss function for differing numbers of sublayers of the same ansatz architecture as the number of qubits increases. The variance was calculated over 500 random initialisations of the variational parameters. The ansatz was created using the generators $[XY,YX,YZ]$ as specified in Figure 6. The red dashed line corresponds to the variance of a normalised average of all the loss functions.
Previously we showed in Equation 40 that the strength of the residual connection determined a lower bound of the probability of success of a given layer. We also noted that the probability of overall success decays exponentially in the number of residual connections $L$. We show in Figure 9 the impact of these effects on the expected number of repeated attempts required to achieve a successful LCU procedure. While the success probability decays exponentially in $L$ it allows the construction of ensembles of size $2^{L}$, which therefore leads to a linear relationship between expected attempts and ensemble size.
Figure 9: The expected repeated attempt to implement one successful LCU procedure (equal to the reciprocal of the overall success probability $\frac{1}{\pi_{S}}$) plotted against the total amount of unitary VQC terms in the ensemble for various fixed residual connection strengths $\beta$. The ensemble will contain $2^{L}$ terms where each term corresponds to a unitary VQC model loss function, where the number of sublayers ranges in increments of $1$ from $\lambda=1$ to $\lambda=2^{L}$. In these numerics we start with an initial state $\rho=\lvert 0\rangle\langle 0\rvert$. At each residual connection layer $l$ we calculate the probability of success of that layer $\pi_{l}$ and then evolve the state by the operator $(I+W^{(\lambda)}_{l})$, where $W^{(\lambda)}_{l}$ consists of $\lambda=2^{l}$ sublayers. In this case we chose each individual sublayer to be a Haar random unitary. The $\beta$ terms correspond to fixing $\beta_{l}=\beta,\forall l$ such that every residual connection has the same strength parameter. Due to the symmetry around $\beta=0.5$ we restricted to plotting $10.5\leq\beta<$ 1 for readability
We also note that in the case that a given residual layer provides an exponentially vanishing contribution to the overall loss, from both its unitary and non-unitary contributions, then this could potentially be used to improve the probability of success scaling of the algorithm by allowing failed implementations of that particular layer to be accepted as successes. In a scenario in which multiple ansätze are being tested simultaneously, in which one may not know which contain barren plateaus or not, it would be possible to ignore any failures for ansätze which do not contribute meaningfully to the loss. This means that the probability of successful implementation would decay exponentially with the number of useful quantum ResNet layers implemented $L^{\prime}$ that provide non-vanishing contributions to the loss, rather than the total number of layers $L$. The implication of this however, is that one could simply discard these unused terms completely from the model and restrict to $L^{\prime}$. In this case an ensemble of $2^{L^{\prime}}$ models could then be manually constructed, which may be able to approximate the quantum ResNet results. Therefore, any benefit from future research in this direction would be focussed on the ability to search through a wide range of ansätze at the same time, and find those that provide meaningful contributions, rather than any quantum advantage inherent in the model performance itself. Although in the worst case settings the cost related to the probabilistic nature of the LCU procedure may render the quantum ResNet infeasible.
Although we show that quantum ResNet frameworks can avoid barren plateaus, it is possible that many common architectures may approach a solution which is vulnerable to classical simulation. This would occur when the quantum ResNet is well approximated by an ensemble of classically simulatable shallow depth terms. Perhaps this situation should be expected, as it has previously been observed that classical deep ResNets can behave as ensembles of shallow networks, where a previous study reported that the gradient of a $110$ layer ResNet was dominated by contributions from paths of depth between $10-34$ layers Veit et al. (2016). We again highlight that the key to further advancements in the quantum ResNet implementation will likely rest on constructing models in which the non-unitary terms do not exponentially vanish and remain hard to simulate classically. This provides an expanded search space compared to strictly unitary models which could be subject to a more in depth characterisation in future work. Overall the question of whether this ensemble-creating property of quantum ResNet can be used effectively therefore remains a question for further research, although the reported success of VQC ensemble models Friedrich and Maziero (2024) may motivate further investigation into the efficient creation of ensembles within a single quantum device.
III Average Pooling Layers
III.1 Quantum Average Pooling Implementation